# Scale Where It Matters: Training-Free Localized Scaling for Diffusion Models

### v2

@@ 4.2 Defect Localization
4.3 Consistency Maintenance
With reliable defect masks from the previous step, the next challenge is how to resample the identified regions without disrupting the rest of the image. Restricting updates only to masked regions often introduces boundary artifacts or semantic drift. To address these issues, LoTTS maintains both spatial and temporal consistency during refinement.
Spatial Consistency.   Resampling in diffusion models typically begins by perturbing the latent representation. If perturbation is applied only within the mask, the noise distribution becomes inconsistent with surrounding regions, creating visible seams. We avoid this by injecting comparable noise into both masked and unmasked areas, which balances noise levels and ensures smooth transitions across boundaries. Formally, given the clean latent $\mathbf{x}_{0}$ and binary mask $\mathbf{M}$, we initialize the perturbed latent at timestep $t_{0}$ as:
$\mathbf{x}_{t_{0}}=\alpha(t_{0})\mathbf{x}_{\mathbf{0}}+\sigma(t_{0})((1-\mathbf{M})\odot\mathbf{z}_{\mathbf{bg}}+\mathbf{M}\odot\mathbf{z}_{\mathbf{mask}})$ (9)
where $\mathbf{z}_{\mathbf{bg}}\!\sim\!\mathcal{N}(0,I)$ and $\mathbf{z}_{\mathbf{mask}}\!\sim\!\mathcal{N}(0,I)$.
Temporal Consistency.   A second challenge is preserving the global semantics of the image. Restarting from pure noise discards structure and forces the model to regenerate the entire scene. Instead, following meng2021sdedit, we resume denoising from an intermediate step $t_{0}$, so that global content is retained while localized corrections remain possible. The masked reverse update is:
$\begin{split}\mathbf{x}_{t-\Delta t}=&\,(1-\mathbf{M})\odot\big(\alpha(t)\mathbf{x}_{0}+\sigma(t)\mathbf{z}\big)\\ &+\mathbf{M}\odot\Big(\mathbf{x}_{t}-\epsilon^{2}\epsilon_{\theta}(\mathbf{x}_{t},t)+\epsilon\mathbf{z}\Big).\end{split}$ (10)
where $\epsilon_{\theta}(\mathbf{x}_{t},t)$ denotes the predicted noise, $\epsilon$ is the per-step noise scale, $N$ is the total number of reverse steps, $\sigma(t)$ is the variance schedule at step $t$, $\mathbf{z}\sim\mathcal{N}(\mathbf{0},\mathbf{I})$, and $\Delta t=t_{0}/N$. This keeps unmasked regions faithful while allowing masked regions to be selectively refined.
Holistic Consistency.   Even after localized refinement, residual inconsistencies may remain around boundaries. To restore smooth transitions, we apply a brief global denoising sweep. This final step enforces full-image coherence and yields consistent, high-fidelity outputs.
5 Experiments and Analysis
Benchmarks.   We evaluate LoTTS on three benchmarks: Pick-a-Pic [pick], DrawBench [saharia2022photorealistic], and COCO2014 [lin2014microsoft]. Performance is measured by four human-aligned metrics, HPSv2 [wu2023human], PickScore [pick], ImageReward (IR) [ir] (also serving as the verifier), and Aesthetic Score (AES) [aes_laion], together with FID [heusel2017gans] and CLIP Score [radford2021learning] on COCO.
Baselines.  Experiments are conducted on SD2.1 [rombach2022high], SDXL [podell2023sdxl], and Flux.1-schnell [flux2024]. SD2.1 and SDXL are diffusion-based, while Flux.1-schnell is a rectified flow-based model. We compare LoTTS with representative sampling and search methods under matched NFE budgets, including vanilla Resampling, Best-of-$N$, and Particle Sampling [kim2025test, singhal2025general].
@@ 5 Experiments and Analysis
Figure 4: Qualitative results on challenging text-to-image prompts. Compared with Resampling, Best-of-NN and Particle Sampling, LoTTS better follows complex prompts. The green borders indicate high-quality generations, red mark lower-quality ones, and brown shows the initial image used as the starting point of localized refinement.
@@ 5.1 Main Results
Quantitative Performance.   Tab.1 shows that LoTTS consistently outperforms Resampling, Best-of-$N$, and Particle Sampling[kim2025test, singhal2025general] under matched NFE budgets across SD2.1, SDXL, and FLUX, achieving clear improvements on both human-aligned metrics (HPS, AES, Pick, IR) and automated metrics. While some baselines occasionally obtain slightly higher AES due to the metric’s preference for stylized, high-contrast renderings, LoTTS focuses on structural and semantic correctness, making small AES fluctuations expected. Importantly, the three benchmarks vary widely in dataset scale and prompt distribution, indicating that LoTTS generalizes well across different evaluation settings. Moreover, the results show that LoTTS is architecture- and sampler-agnostic: it works well with both U-Net-based and Transformer-based diffusion model and suitable for both SDE- and ODE-based samplers.
Qualitative Comparison.   Figure 4 shows that LoTTS produces cleaner and more coherent images across challenging prompts. The w/o Scaling results often contain severe artifacts. For example, muddy textures, incomplete patterns, distorted metal parts, and missing wolf for those prompts. Best-of-$N$ and Particle Sampling substantially improve quality but still suffer from issues such as inconsistent lighting, oversaturated colors, and local geometric distortions. In contrast, LoTTS performs targeted local refinements on the initial generation (with brown border), improving semantic details while preserving the global layout. These observations demonstrate that LoTTS more effectively improves visual fidelity and stability than existing TTS baselines.
Figure 5: Localized Refinement Results on SD2.1. LoTTS corrects diverse localized artifacts, such as distorted object geometry, duplicated fingers, texture noise, and warped text. Circles highlight corrected regions.
Local Refinement Capability.   Fig. 5 illustrates the different types of localized artifacts that LoTTS effectively corrects. The cow example shows a distorted or missing head; the hand example exhibits duplicated or fused fingers; the bicycle contains wavy, noisy textures on the spokes; and the airplane displays warped, unreadable text on the fuselage. Across these diverse cases, LoTTS reliably refines the corrupted regions while preserving surrounding content, demonstrating both the effectiveness of localized resampling and its applicability to a broad range of artifact types. White and Red circles both mark the corrected areas.
More Results.   More qualitative results and failure cases provided in Appendix C and F.
@@ 5.2 Ablations
Mask Generation.   Fig.6 visualizes the intermediate outputs of our mask generation pipeline, showing how each stage contributes to coherent defect localization. Remarkably, despite being unsupervised, our final mask qualitatively appears more precise and semantically aligned than the supervised defect-detection model [li2024g]. As summarized in Fig.7, removing any component – whether the discriminative difference map $\mathrm{CA}_{\mathrm{diff}}^{\prime}$, the propagated original attention $\mathrm{CA}_{\mathrm{orig}}^{\prime}$, or the context-aware propagation – consistently degrades all four metrics. The variant without $\mathrm{CA}_{\mathrm{orig}}^{\prime}$ performs the worst overall, indicating the importance of the strong semantic foreground prior. The no-mask and random-mask baselines also underperform, further confirming the need for accurate, structure-aware masks.
Figure 6: Visualization of Cross-Attention–based Defect Localization. From the original ($\mathrm{CA}_{\mathrm{orig}}$), positive ($\mathrm{CA}_{\mathrm{pos}}$), and negative ($\mathrm{CA}_{\mathrm{neg}}$) prompts, LoTTS derives discriminative differences ($\mathrm{CA}_{\mathrm{diff}}^{\prime}$) and combines them with propagated original attention ($\mathrm{CA}_{\mathrm{orig}}^{\prime}$) to produce a coherent defect mask ($\mathrm{CA}_{\mathrm{diff}}^{\prime}+\lambda\mathrm{CA}_{\mathrm{orig}}^{\prime}$) that highlights low-quality regions. The Reference shows the masks from a suboptimal supervised model [li2024g].
Figure 7: Ablation of Mask Generation for LoTTS. Mask generation ablations show each component contributes to generation performance.
Quality Prompt Design.   As shown in Tab. 2, different auxiliary prompts (e.g., “a good/bad image of {p}”) yield similar results, indicating that LoTTS is robust to prompt phrasing. These prompts serve only to elicit contrastive attention patterns – “good” prompts highlight regions the model considers well-formed, while “bad” prompts emphasize areas linked to low quality. The consistent performance across phrasing suggests that LoTTS relies on this quality contrast rather than the specific wording, and that diffusion models encode a stable notion of visual quality that our LoTTS can reliably exploit.
Table 2: Ablation of Prompt Design for LoTTS. “Good/Bad” uses the prompts “a good/bad image of {p}”. “High/Low-quality” uses “a high/low-quality image of {p}”.
@@
Good/Bad 6.29 28.62 22.21 1.140
High/Low-quality 6.29 28.69 22.29 1.137
Figure 8: Parameter sensitivity analysis of LoTTS. Varying localized refinement iterations ($k$), resample area ratio ($r$), and denoising start step ($t_{0}$) shows that LoTTS maintains stable improvements across HPS, AES, Pick, and IR metrics.
Number of Refinements ($k$).   As shown in Fig. 8, performance peaks at $k=2$. Due to the stochastic nature of diffusion sampling, a single refinement may miss the correct reconstruction, while $2$ attempts provide enough diversity to reliably fix most artifacts. Beyond this point, additional refinements rarely offer new solutions, suggesting that the effective correction space is reached within just a few attempts.
Mask Area Ratio ($r$).   Moderate mask ratio r=50r=50% works best. Small masks tend to miss parts of the defects, while large masks begin to overwrite clean regions and behave like full-image resampling. Overall, LoTTS is fairly robust to the choice of $r$, showing stable performance across a wide range of mask sizes.
Noise Injection Step ($t_{0}$).   The optimal noise injection step depends on the model’s generative ability. Injecting noise too early disrupts too much original information, while injecting too late provides too little flexibility to correct defects, so $t_{0}$ must balance structural preservation with refinability. For diffusion models like SD2.1 and SDXL, mid-range steps work well, whereas FLUX converges very early and therefore benefits from injecting noise closer to the noise-dominated steps.
Scaling Efficiency.   As shown in Fig. 9, LoTTS reaches Best-of-$N$ performance with far fewer samples, yielding $2.8\times$$3.3\times$ speedups on SD2.1 and SDXL and up to $4\times$ on FLUX. The larger speedup on stronger models shows that even high-quality generators still leave local defects that LoTTS can efficiently fix, positioning localized TTS as an increasingly essential tool in the next generation of large generative models.
Figure 9: Scaling Comparison on LoTTS vs. Best-of-NN Sampling.
More Ablations.   Additional ablation and parameter sensitivity results are included in Appendix D and E.
@@ 5.3 Discussion
As shown in Fig.10, LoTTS occasionally produces failures. We think LoTTS failures mainly fall into three categories. First, imperfect masks: extremely subtle defects or high-frequency regions may cause missed or unstable attention signals. This challenge is not unique to our training-free LoTTS – supervised defect-detection models exhibit similar limitations – but importantly, our Appendix G shows that localized TTS can outperform global resampling even with imperfect masks. Second, global defects: issues involving perspective, spatial relations, or global semantics cannot be fixed through local resampling alone, pointing toward future work on adaptive local–global TTS strategies. Third, initial images at quality extremes: low-quality images may be too fragile to refine locally, while near-perfect images offer little room for improvement, suggesting that early-stopping mechanisms may further improve LoTTS. Overall, these cases point to natural opportunities for extending localized TTS, and LoTTS marks an early step toward more adaptive scaling methods.
@@ 6 Conclusion
We proposed LoTTS, a training-free framework that extends test-time scaling from global resampling to localized refinement. Unlike conventional methods that apply sampling uniformly, LoTTS leverages defect-aware masks and consistency constraints to focus computation where it matters most. This design not only improves image quality and efficiency but also shows that scaling can be made adaptive to the spatial heterogeneity of generative outputs. Beyond diffusion models, the principle of region-aware scaling provides a general perspective for developing more efficient and controllable inference strategies in generative AI. Our analysis further establishes theoretical conditions under which localized scaling provably outperforms global sampling, and we believe LoTTS opens up promising directions for integrating fine-grained control into test-time algorithms across diverse generative architectures.
[…]
Supplementary Material
AImplementation Details...........................................................................................................A
BAlgorithmic Details...........................................................................................................B
  B.1Notation and Preliminaries...........................................................................................................B.1
  B.2Defect Localization Procedure...........................................................................................................B.2
  B.3Localized Resampling Procedure...........................................................................................................B.3
  B.4Overall Algorithm of LoTTS...........................................................................................................B.4
CQualitative Results...........................................................................................................C
  C.1Mask Generation Examples...........................................................................................................C.1
  C.2Localized Refinement Examples...........................................................................................................C.2
DAblations...........................................................................................................D
  D.1Mask Generation...........................................................................................................D.1
  D.2Prompt Design...........................................................................................................D.2
EParameter Analysis...........................................................................................................E
  E.1Hyperparameter Sensitivity...........................................................................................................E.1
  E.2Scaling Comparison with Best-of-$N$...........................................................................................................E.2
FDiscussion...........................................................................................................F
  F.1Failure Analysis...........................................................................................................F.1
  F.2Future Directions...........................................................................................................F.2
GTheoretical Analysis...........................................................................................................G
  G.1Problem Setup and Preliminaries...........................................................................................................G.1
  G.2Main Theorem...........................................................................................................G.2
  G.3Derived Corollaries...........................................................................................................G.3
  G.4Practical Q&A...........................................................................................................G.4
@@ Appendix A Implementation Details
Hardware and software.   All experiments were run on NVIDIA A800 (80GB) GPUs; most runs use a single GPU unless otherwise stated. Mixed precision is enabled via torch.cuda.amp.autocast (fp16) unless the metric implementation requires fp32.
@@
Default sampler and schedule.   For SD2.1, we use SDE-DPM-Solver++ sampling with 50 NFEs, and CFG scale =7.5=7.5. For SDXL, we use SDE-DPM-Solver++ sampling with 30 NFEs, and CFG scale =5.5=$5.$ For FLUX, we use modified SDE-DPM-Solver++ sampling with 10 NFEs, and no CFG. For LoTTS, we inject noise to reach an intermediate step $t_{0}$ (Sec. 4), then perform local masked refinement followed by a short global integration phase. We keep the original prompt and CFG weight unchanged during refinement.
Attention hook and mask resolution.   For SD2.1, we hook U-Net cross-attention tensors at the $16{\times}16$ spatial resolution blocks (for $512{\times}512$ images). For SDXL, we hook U-Net cross-attention tensors at the $64{\times}64$ spatial resolution blocks (for $1024{\times}1024$ images). For FLUX, we extract Transformer cross-attention tensors from the $64{\times}64$ resolution of the last 10 blocks (for $1024{\times}1024$ images), and generate attention maps following the method of [helbling2025conceptattention]. The attention maps are averaged across heads, mean-pooled across the feature dimension, and then averaged across tokens to obtain a single spatial map per prompt type (pos/neg/origin). Self-attention maps are normalized to $[0,1]$. The quality map $P$ uses equation 7 in the paper with $\lambda=0.5$ by default. We compute attention and the final mask at t=0t{=}0 (last step), where maps and image latents are already stable. Masks are kept in latent resolution during resampling; only for visualization we bilinearly upsample to pixel space.
Datasets.   For Pick-a-Pic, we follow the official prompt pool for preference-oriented evaluation. For DrawBench, we use all categories and report aggregate metrics. For COCO2014, we use the widely-used COCO2014-30k.
Metrics.   We report HPSv2, PickScore, ImageReward, and AES using the authors’ official weights and preprocessing. CLIP score and FID are computed with standard settings: CLIP uses ViT-L/14 unless otherwise stated; FID uses Inception-V3 features, $2048$-D activations. ImageReward (IR) score is used as the global selection score for Best-of-$N$, LoTTS, and other search-based baselines.
Reproduction checklist.   (1) Use the same checkpoint and VAE; (2) match resolution, sampler, steps, CFG; (3) compute the mask at t=0t{=}0 with the same prompt templates; (4) keep $\lambda$, $r$, $t_{0}$, $T^{\prime}$, $k$ as in Tab. 1; (5) fix seeds and determinism flags; (6) run the same metric code and preprocessing.
@@
Sampler SDE-DPM- Solver++ SDE-DPM- Solver++ Modified SDE- DPM-Solver++
@@
[…]
Table 1: Default LoTTS hyperparameters. Transposed single-column version with wrapped sampler names.
Appendix B Algorithmic Details
B.1 Notation and Preliminaries
@@
$\mathcal{L}$ set of U-Net layers used to extract cross-attention maps
@@
$S$ number of global seeds at DFS depth 1
@@
NN number of refinement steps (also for $t_{g}\!\to\!0$ discretization)
@@
$\mathrm{CrossAttn}$ cross-attention maps from U-Net layers $\mathcal{L}$ during a forward pass with latent $x$ and prompt $p$
Table 2: Unified notation for Algorithms 1, 2, and 3.

### v3

@@ 5 Experiments and Analysis
Benchmarks. We evaluate LoTTS on three benchmarks: Pick-a-Pic (pick), DrawBench (saharia2022photorealistic), and COCO2014 (lin2014microsoft). Performance is measured by four human-aligned metrics, HPS(v2) (wu2023human), PickScore (pick), ImageReward (IR) (ir), and Aesthetic Score (AES) (aes_laion), together with FID (heusel2017gans) and CLIP Score (radford2021learning) on COCO. Following prior work (kim2025test; singhal2025general), all methods use the same verifier (IR) to select the best candidate, ensuring a fair comparison.
Baselines. Experiments are conducted on SD2.1 (rombach2022high), SDXL (podell2023sdxl), and Flux.1-schnell (flux2024). SD2.1 and SDXL are diffusion-based, while Flux.1-schnell is a rectified flow-based model. We compare LoTTS with representative sampling and search methods under matched NFE budgets, including vanilla Resampling (standard sampling with a single random seed and no verifier selection), Best-of-$N$, and Particle Sampling (kim2025test; singhal2025general).
@@
Figure 6: Localized Refinement on SD2.1. LoTTS corrects diverse localized artifacts (e.g. distorted hands and faces). Heatmaps show per-region artifact scores, and binary masks indicate the regions selected for refinement.
@@ 5.1 Main Results
Quantitative Performance. Tab. 1 shows that LoTTS consistently outperforms Resampling, Best-of-$N$, and Particle Sampling (kim2025test; singhal2025general) under matched NFE budgets across SD2.1, SDXL, and FLUX, achieving clear improvements on both human-aligned metrics (HPS, AES, Pick, IR) and automated metrics. Notably, the gains extend beyond the verifier metric (IR) to independent metrics such as HPS, PickScore, FID, and CLIP, confirming genuine quality improvement rather than verifier overfitting. The consistent gains across three benchmarks and three architectures (U-Net-based SD2.1/SDXL and Transformer-based FLUX) confirm that LoTTS generalizes well and is architecture-agnostic.
Human Evaluation. We conduct a blind pairwise preference study on SD2.1 with 100 prompts. Five human annotators independently choose the preferred image from each pair without knowing the method identity, yielding 500 total judgments. LoTTS is preferred over Best-of-$N$ in 86% of cases, consistent with the automatic metric trends.
Qualitative Comparison. Figure 5 shows that unscaled results often contain severe artifacts, while Best-of-$N$ and Particle Sampling still exhibit issues such as inconsistent lighting and local distortions. Starting from the initial sample (brown border), LoTTS performs targeted local refinements while preserving the global layout, producing more stable and higher-fidelity images.
Local Refinement Capability. Fig. 6 shows several examples of localized refinement by LoTTS. The flamingo case corrects a duplicated bird head near the legs, while the kitten example introduces the missing yarn mentioned in the prompt. For the wizard tower and floating islands, refinement mainly enhances structural details and visual clarity. Heatmaps highlight artifact-prone regions, and binary masks indicate the areas selected for localized resampling, leading to consistent improvements in ImageReward (IR).
More Results. Additional qualitative results on SDXL and FLUX, as well as failure cases, are provided in Appendix C and F.
@@ 5.2 Ablations
Mask Generation. As summarized in Fig. 8(1), removing any component – whether the discriminative difference map $\mathrm{CA}_{\mathrm{diff}}^{\prime}$, the propagated original attention $\mathrm{CA}_{\mathrm{orig}}^{\prime}$, or the context-aware propagation – consistently degrades all four metrics. The variant without $\mathrm{CA}_{\mathrm{orig}}^{\prime}$ performs the worst overall, indicating the importance of the strong semantic foreground prior. The no-mask and random-mask baselines also underperform, further confirming the need for accurate, structure-aware masks.
External Validation of Mask Quality. To verify that our masks reflect genuine quality signals, we compare them with pixel-level quality heatmaps from MUSIQ (ke2021musiq), an independent NR-IQA model, computed via sliding windows (patch 64, stride 32). As shown in Fig. 7, both methods consistently highlight similar low-quality regions, despite ours being entirely training-free. This confirms that diffusion cross-attention encodes meaningful perceptual quality information that LoTTS can exploit without any additional training.
Figure 7: Comparison with MUSIQ (ke2021musiq) quality maps. Each triplet shows the original image, our attention heatmap, and the MUSIQ heatmap (red = lower quality). Both methods consistently highlight similar low-quality regions.
Figure 8: (1) Ablation of Mask Generation. Removing any component consistently degrades all four metrics. (2) Prompt Design. Radar chart comparing two prompt strategies: “Good/Bad” (“a good/bad image of {p}”) and “High/Low-Quality” (“a high/low-quality image of {p}”). The nearly overlapping profiles show that LoTTS is robust to prompt phrasing.
Quality Prompt Design. As shown in Fig. 8(2), different auxiliary prompts (e.g. “a good/bad image of {p}”) yield nearly identical metric profiles, indicating that LoTTS is robust to prompt phrasing and relies on quality contrast rather than specific wording.
Figure 9: (1) Parameter sensitivity analysis of LoTTS. Varying localized refinement iterations ($k$), resample area ratio ($r$), and denoising start step ($t_{0}$) shows that LoTTS maintains stable improvements across HPS and AES metrics. (2) Scaling Comparison: LoTTS vs. Best-of-$N$ Sampling. LoTTS reaches the same HPS with far fewer samples, yielding significant speedups.
Table 2: Wall-clock latency comparison. Per-image latency (seconds) and total time to reach comparable quality. LoTTS achieves similar per-image cost but significantly lower total runtime.
SD2.1 SDXL FLUX
Method Per-img Total Per-img Total Per-img Total
Best-of-NN 0.613 13.80 0.294 7.050 0.154 3.888
Particle Sampling 0.616 8.31 0.298 4.760 0.157 2.750
LoTTS (Ours) 0.618 5.01 0.298 2.16 0.158 1.04
@@
Parameter Sensitivity ($k$, $r$, $t_{0}$). As shown in Fig. 9(1), LoTTS is robust across a wide range of hyperparameters. Performance peaks at k=2k{=}2 refinements, as two attempts provide sufficient stochastic diversity to fix most artifacts. A moderate mask ratio $r{=}50\%$ works best: smaller masks miss defects while larger ones overwrite clean regions. The noise injection step $t_{0}$ balances preservation and refinability – mid-range steps suit SD2.1/SDXL, while FLUX benefits from earlier injection due to its rapid convergence.
Scaling Efficiency. As shown in Fig. 9(2), LoTTS reaches the same HPS as Best-of-$N$ with $2.8\times$$3.3\times$ fewer samples on SD2.1 and SDXL, and up to $4\times$ fewer on FLUX, with larger savings on stronger models.
Wall-clock Latency. Tab. 2 translates the above sample savings into wall-clock time: LoTTS achieves comparable per-image latency while requiring far fewer samples to reach the target quality, resulting in $2.75\times$$3.76\times$ lower total runtime.
More Ablations. Further ablations on $\lambda$ and the consistency components, along with extended parameter analyses, are provided in Appendix D and E.
@@ 5.3 Discussion
As shown in Fig.10, LoTTS occasionally produces failures. We observe that LoTTS failures mainly fall into three categories. First, imperfect masks: extremely subtle defects or high-frequency regions may cause missed or unstable attention signals. This challenge is not unique to our training-free LoTTS – supervised defect-detection models exhibit similar limitations – but importantly, our Appendix G shows that localized TTS can outperform global resampling even with imperfect masks. Second, global defects: issues involving perspective, spatial relations, or global semantics cannot be fixed through local resampling alone, pointing toward future work on adaptive local–global TTS strategies. Third, initial images at quality extremes: low-quality images may be too fragile to refine locally, while near-perfect images offer little room for improvement, suggesting that early-stopping mechanisms may further improve LoTTS. Overall, these cases point to natural opportunities for extending localized TTS, and LoTTS marks an early step toward more adaptive scaling methods.
@@ 6 Conclusion
We proposed LoTTS, a training-free framework that extends test-time scaling from global resampling to localized refinement. Unlike conventional methods that apply sampling uniformly, LoTTS leverages defect-aware masks and consistency constraints to focus computation where it matters most. This design not only improves image quality and efficiency but also shows that scaling can be made adaptive to the spatial heterogeneity of generative outputs. Beyond the specific models tested, we believe the principle of region-aware scaling can benefit a broader class of generative architectures. Our analysis further establishes that localized scaling provably outperforms global sampling under mild conditions, and we hope LoTTS opens up promising directions for integrating fine-grained, spatially adaptive control into test-time algorithms.
@@ Appendix A Implementation Details
Hardware and software. All experiments were run on NVIDIA A100 (80GB) GPUs; most runs use a single GPU unless otherwise stated. Mixed precision is enabled via torch.cuda.amp.autocast (fp16) unless the metric implementation requires fp32.
@@
Default sampler and schedule. For SD2.1, we use SDE-DPM-Solver++++ sampling with 50 NFEs, and CFG scale =7.5=7.5. For SDXL, we use SDE-DPM-Solver++++ sampling with 30 NFEs, and CFG scale =5.5=$5.$ For FLUX, we use modified SDE-DPM-Solver++++ sampling with 10 NFEs, and no CFG. For LoTTS, we inject noise to reach an intermediate step $t_{0}$ (Sec. 4), then perform local masked refinement followed by a short global integration phase. We keep the original prompt and CFG weight unchanged during refinement.
Table 3: Default LoTTS hyperparameters.
@@
Sampler SDE-DPM- Solver++++ SDE-DPM- Solver++++ Modified SDE- DPM-Solver++++
@@
$r\;(\%)$ 50 50 50
Attention hook and mask resolution. For SD2.1, we hook U-Net cross-attention tensors at the $16{\times}16$ spatial resolution blocks (for $512{\times}512$ images). For SDXL, we hook U-Net cross-attention tensors at the $64{\times}64$ spatial resolution blocks (for $1024{\times}1024$ images). For FLUX, we extract Transformer cross-attention tensors from the $64{\times}64$ resolution of the last 10 blocks (for $1024{\times}1024$ images), and generate attention maps following the method of (helbling2025conceptattention). The attention maps are averaged across heads, mean-pooled across the feature dimension, and then averaged across tokens to obtain a single spatial map per prompt type (pos/neg/origin). Self-attention maps are normalized to $[0,1]$. The quality map $P$ is computed via Eq. equation 7 with $\lambda=0.5$ by default. We compute attention and the final mask at t=0t{=}0 (last step), where maps and image latents are already stable. Masks are kept in latent resolution during resampling; only for visualization we bilinearly upsample to pixel space.
Datasets. For Pick-a-Pic, we follow the official prompt pool for preference-oriented evaluation. For DrawBench, we use all categories and report aggregate metrics. For COCO2014, we use the standard COCO2014-30k split.
Metrics. We report HPSv2, PickScore, ImageReward, and AES using the authors’ official weights and preprocessing. CLIP score and FID are computed with standard settings: CLIP uses ViT-L/14 unless otherwise stated; FID uses Inception-V3 features with $2048$-D activations. ImageReward (IR) score is used as the global selection score for Best-of-$N$, LoTTS, and other search-based baselines.
Reproduction checklist. (1) Use the same checkpoint and VAE; (2) match resolution, sampler, steps, CFG; (3) compute the mask at t=0t{=}0 with the same prompt templates; (4) keep $\lambda$, $r$, $t_{0}$, $t_{g}$, $k$ as in Tab. 3; (5) fix seeds and determinism flags; (6) run the same metric code and preprocessing.
Human Evaluation Protocol. We conduct a blind pairwise preference study as described in Sec. 5. Fig. 11 shows the annotation interface: each annotator views two images side-by-side without method labels and selects the preferred one. Five annotators independently evaluated 100 prompts, yielding 500 judgments total.
Table 4: Unified notation for Algorithms 1, 2, and 3.
@@
$\mathcal{L}$ set of backbone layers used to extract cross-attention maps
@@
$S$ number of global seeds at DFS depth 0 (root)
@@
$\Delta t$ scheduler step size used for discretization
@@
$\mathrm{CrossAttn}$ cross-attention maps from backbone layers $\mathcal{L}$ during a forward pass with latent $x$ and prompt $p$
Appendix B Algorithmic Details
B.1 Notation and Preliminaries